# Learning large softmax mixtures with warm start EM

### v1

@@ B.3 Proofs of Lemmas 2, 3 and 4
In view of (B.3.1) and (B.3.1) and the fact that ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},r)$ implies ${\bm{\theta}}_{k},{\bm{\theta}}_{k}^{\prime}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$ hence ${\bm{\theta}}_{k,u}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$ for any $u\in[0,1]$, (78) and (79) follow if the event
$\mathcal{E}_{H}:=\bigcap_{k=1}^{K}\left\{(1-2c_{0})\underline{\sigma}^{2}\leq\lambda_{L}(H_{{\bm{\theta}}_{k}})\leq\lambda_{1}(H_{{\bm{\theta}}_{k}})\leq(1+2c_{0})\overline{\sigma}^{2},~{}\forall{\bm{\theta}}_{k}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0})\right\}$ (95)
holds with probability
$\mathbb{P}\left\{\mathcal{E}_{H}\right\}\geq 1-\mathcal{O}(Kp^{1-s}).$ (96)
To prove (96), write $H_{{\bm{\theta}}^{*}}^{1/2}$ as the matrix square root of $H_{{\bm{\theta}}^{*}}=\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}^{*}}}\mathbf{X}$ for any ${\bm{\theta}}^{*}\in\{{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}\}$. Let ${\bm{\theta}}\in{\mathbb{B}}({\bm{\theta}}^{*},\delta_{0}/\overline{\sigma})$ be arbitrary. We first bound
$=\sup_{v\in\mathbb{S}^{L-1}}\left(R_{1}(v)+R_{2}(v)\right)$
with
$R_{1}(v):=\sum_{j=1}^{p}(X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}\left(A_{{\bm{\theta}}^{*}}(X_{j})-A_{{\bm{\theta}}}(X_{j}))\right)$
and
$R_{2}(v)=\left|\sum_{j=1}^{p}X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(X_{j})+A_{{\bm{\theta}}}(X_{j})\right)\right|~{}\left|\sum_{j=1}^{p}X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(X_{j})-A_{{\bm{\theta}}}(X_{j})\right)\right|.$
$\leq 3c_{0}\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}$ by (88)
[…]
Next, we observe that
$R_{2}(v)$ $\leq 2\left|\sum_{j=1}^{p}(X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)A_{{\bm{\theta}}^{*}}(X_{j})\right|\left|\sum_{j=1}^{p}(X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(X_{j})-A_{{\bm{\theta}}}(X_{j}))\right|$
$+\left|\sum_{j=1}^{p}(X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(X_{j})-A_{{\bm{\theta}}}(X_{j}))\right|^{2}.$
By repeating the arguments in (97), we find that
$\leq 3c_{0}\left(\sum_{j=1}^{p}(X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}A_{{\bm{\theta}}^{*}}(X_{j})\right)^{1/2}\left(\sum_{j=1}^{p}A_{{\bm{\theta}}^{*}}(X_{j})\right)^{1/2}$
$\leq 3c_{0}\sqrt{\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}}$
$=3c_{0}$ (99)
so that
$\sup_{v\in\mathbb{S}^{L-1}}R_{2}(v)$ $\leq 6c_{0}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}\left|X_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\right|A_{{\bm{\theta}}^{*}}(X_{j})+9c_{0}^{2}\leq 3(2c_{0}+3c_{0}^{2}).$
$\|H_{{\bm{\theta}}^{*}}^{-1/2}(H_{\bm{\theta}}-H_{{\bm{\theta}}^{*}})H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}\leq 9c_{0}(1+c_{0})$
Together with Weyl’s inequality, this gives
$\left|1-\lambda_{\ell}(H_{{\bm{\theta}}^{*}}^{-1/2}H_{\bm{\theta}}H_{{\bm{\theta}}^{*}}^{-1/2})\right|\leq 9c_{0}(1+c_{0}),\quad\forall~{}1\leq\ell\leq L.$
In particular,
$(1-9c_{0}-9c_{0}^{2})\lambda_{L}(H_{{\bm{\theta}}^{*}})\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq\lambda_{1}(H_{{\bm{\theta}}^{*}})(1+9c_{0}+9c_{0}^{2}).$
Invoking Lemma 13 with ${\bm{\theta}}^{*}={\bm{\theta}}_{k}^{*}$, ${\bm{\theta}}={\bm{\theta}}_{k}$ and taking the union bounds over $k\in[K]$ gives (96), thereby completing the proof. ∎
B.3.2 Proof of Lemma 3: the gradient smoothness of the surrogate function $Q$ and the Lipschitz continuity of $M_{\alpha}$
@@
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ be arbitrary. We first recall that along almost all sample sequences $X_{1},X_{2},\cdots$, given $(X_{1},\ldots,X_{p})$, as $p\to\infty$, the results in Lemma 6 and Lemma 7 hold.
We first prove (82). From (51), we argue
@@
Here, we write ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{*}$. Further we denote
@@
After we invoke Lemmas 6 and 7, we find that
@@
Collecting (103), (104) as well as (102) yields that, with the same probability,
@@
Next we prove (81). By definition (49), we can split, for any $k\in[K]$,
@@
$\max_{k\in[K]}|M_{k}({{\bm{\omega}}})-M_{k}({{\bm{\omega}}^{*}})\|$ $\lesssim\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}+\left(1+\overline{\sigma}\Delta\right)\overline{\sigma}\max_{a\in[K]}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}\right)$
@@
B.3.3 Proof of Lemma 4: concentration inequality of the EM-updates within the specified neighborhood
@@
${\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta)=\left\{{\bm{\omega}}\in\Omega:\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\leq\delta_{0}\underline{\alpha},~{}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma},~{}\forall~{}k\in[K]\right\}.$
For any given $\epsilon_{1}\in(0,\underline{\alpha}/4]$ and $\epsilon_{2}\in(0,\delta_{0}/\overline{\sigma}]$, let $\mathcal{N}_{\epsilon_{1}}$ be an $\epsilon_{1}$-covering set (in $\ell_{\infty}$-norm) of $\Delta_{K}$ and $\mathcal{N}_{\epsilon_{2}}$ be the $\epsilon_{2}$-net (in $\ell_{2}$-norm) of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Then, for any $k\in[K]$, $\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{k}^{*}):=\left\{{\bm{\theta}}+{\bm{\theta}}_{k}^{*}:{\bm{\theta}}\in\mathcal{N}_{\epsilon_{2}}\right\}$ is the $\epsilon_{2}$-net of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Consider the set
$\mathcal{N}_{\epsilon_{1},\epsilon_{2}}=\mathcal{N}_{\epsilon_{1}}\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{1}^{*})\otimes\cdots\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{K}^{*}).$
@@
$\|{\bm{\alpha}}^{*}-{\bm{\alpha}}^{\prime}\|_{\infty}\leq\epsilon_{1}+\delta_{0}\underline{\alpha},\quad\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k}^{\prime}\|_{2}\leq\delta_{0}\overline{\sigma}.$ (109)
@@
$|\mathcal{N}_{\epsilon_{1},\epsilon_{2}}|\leq|\mathcal{N}_{\epsilon_{1}}(\alpha^{*})||\mathcal{N}_{\epsilon_{2}}|\leq\left(5\over\epsilon_{1}\right)^{K-1}\left(3\delta_{0}\over\overline{\sigma}\epsilon_{2}\right)^{L}.$ (110)
Since for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, there exists ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (108) – (109) such that
$\begin{split}\max_{k\in[K]}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|&\leq\max_{k\in[K]}|\widehat{M}_{k}({\bm{\omega}}^{\prime})-M_{k}({\bm{\omega}}^{\prime})|\\ &\quad+\max_{k\in[K]}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})|\end{split}$ (111)
We first bound the second term.
$\begin{split}&\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})\right|\\ &=\max_{k\in[K]}\left|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(X_{j}))\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{{\bm{\omega}}}(X_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\right)\right|\\ &\leq\|\widehat{\pi}-\pi_{{\bm{\omega}}^{*}}\|_{1}\max_{k\in[K],j\in[p]}\left|{\alpha_{k}A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{{\bm{\omega}}}(X_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\right|\\ &\leq 2\max_{k\in[K],j\in[p]}\left(|\alpha_{k}-\alpha_{k}^{\prime}|{A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{{\bm{\omega}}}(X_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(X_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}{|\pi_{{\bm{\omega}}}(X_{j})-\pi_{{\bm{\omega}}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}\right)\\ &\leq 2\max_{k\in[K],j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{{\bm{\omega}}}(X_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(X_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}{|\pi_{{\bm{\omega}}}(X_{j})-\pi_{{\bm{\omega}}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}\right)\\ &\leq 2\max_{k\in[K],j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(X_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}+{|\pi_{{\bm{\omega}}}(X_{j})-\pi_{{\bm{\omega}}^{\prime}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}\right).\end{split}$ (112)
@@
Hence, we can invoke Lemma 5, and obtain
@@
$\quad\lesssim\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}$ (114)
@@
with $[h_{k}]_{j}:=\alpha_{k}A_{{\bm{\theta}}_{k}}(X_{j})/\pi_{{\bm{\omega}}}(X_{j})$ for all $j\in[p]$ and $E_{1},\ldots,E_{N}$ are samples of Multinomial$(N;\pi_{{\bm{\omega}}^{*}})$. Note that
@@
$\epsilon_{1}=\underline{\alpha}\left({1\over 4}\wedge{(L+K)\over N}\right),\qquad\epsilon_{2}={1\over\overline{\sigma}}\left(\delta_{0}\wedge{\overline{\sigma}\over\|\mathbf{X}\|_{\infty,2}}{(L+K)\over N}\right)$
as well as $t=C(L+K)\log(N)$.
@@
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},r)}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}$
$\leq\max_{{\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})\|_{2}$
@@
Pick any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},r)$ and ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (108). We bound from above
@@
where $\mathcal{N}_{L}(1/2)$ is the $(1/2)$-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/2)|\leq 5^{L}$ (see, for instance, Vershynin (2018)). We work on the event $\mathcal{E}_{H}$ in (95) satisfying (96). Fix any ${\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ and $v\in\mathcal{N}_{L}(1/2)$. Observe that
@@
$\leq\alpha_{k}\max_{j\in[p]}\left(1+{|\pi_{{\bm{\omega}}}(X_{j})-\pi_{{\bm{\omega}}^{*}}(X_{j})|\over\pi_{{\bm{\omega}}}(X_{j})}\right)\alpha_{k}\lambda_{1}\left(\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}_{k}}}\mathbf{X}\right)$
$\overset{(i)}{\lesssim}\alpha_{k}\left(1+\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}+\delta_{0}\right)\overline{\sigma}^{2}$
$\overset{(ii)}{\lesssim}\alpha_{k}\overline{\sigma}^{2}$
where the step $(i)$ uses (91), (95) and (109) while the step $(ii)$ is due to (88), $\epsilon_{2}\leq{\delta_{0}/\overline{\sigma}}$ and $\epsilon_{1}\leq\underline{\alpha}/4$. By further noticing

### v2

@@ B.2.1 Technical lemmas used to prove Lemmas 4, 5 and 6
In view of (B.2.2) and (B.2.2) and the fact that ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ implies ${\bm{\theta}}_{k},{\bm{\theta}}_{k}^{\prime}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$, the Euclidean ball around ${\bm{\theta}}_{k}^{*}$ with radius $\delta_{0}/\overline{\sigma}$, so that ${\bm{\theta}}_{k,u}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$ for any $u\in[0,1]$, (65) and (66) follow by invoking Lemma 8 with ${\bm{\theta}}^{*}={\bm{\theta}}_{k}^{*}$ and ${\bm{\theta}}_{k}={\bm{\theta}}_{k,u}$ for all $k\in[K]$. The proof is complete. ∎
B.2.3 Proof of Lemma 5: the gradient smoothness of the surrogate function $Q$ and the Lipschitz continuity of $M_{\alpha}$
@@
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ be arbitrary. We first prove (69). From (14), we argue
@@
Here, we recall ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{*}$. Further we denote
@@
After we invoke Lemmas 9 and 10, we find that
@@
Collecting (86), (87) as well as (85) yields that
@@
Next we prove (68). By definition in (12), we can split, for any $k\in[K]$,
@@
$\max_{k\in[K]}|M_{k}({{\bm{\omega}}})-M_{k}({{\bm{\omega}}^{*}})|$ $\lesssim\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}+\left(1+\overline{\sigma}\Delta\right)\overline{\sigma}\max_{a\in[K]}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}\right)$
@@
B.2.4 Proof of Lemma 6: concentration inequality of the EM-updates within the specified neighborhood
@@
${\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})=\left\{{\bm{\omega}}:\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\leq\delta_{0}\underline{\alpha},~\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma},~\forall~k\in[K]\right\}.$
For any given $\epsilon_{1}\in(0,\underline{\alpha}/4]$ and $\epsilon_{2}\in(0,\delta_{0}/\overline{\sigma}]$, let $\mathcal{N}_{\epsilon_{1}}(\Delta^{K})$ be an $\epsilon_{1}$-covering set (in $\ell_{\infty}$-norm) of $\Delta^{K}$ and $\mathcal{N}_{\epsilon_{2}}$ be the $\epsilon_{2}$-net (in $\ell_{2}$-norm) of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Then, for any $k\in[K]$, $\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{k}^{*}):=\left\{{\bm{\theta}}+{\bm{\theta}}_{k}^{*}:{\bm{\theta}}\in\mathcal{N}_{\epsilon_{2}}\right\}$ is the $\epsilon_{2}$-net of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Consider the set
$\mathcal{N}_{\epsilon_{1},\epsilon_{2}}=\mathcal{N}_{\epsilon_{1}}(\Delta^{K})\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{1}^{*})\otimes\cdots\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{K}^{*}).$
@@
$\|{\bm{\alpha}}^{*}-{\bm{\alpha}}^{\prime}\|_{\infty}\leq\epsilon_{1}+\delta_{0}\underline{\alpha},\quad\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k}^{\prime}\|_{2}\leq\delta_{0}/\overline{\sigma}.$ (92)
@@
$|\mathcal{N}_{\epsilon_{1},\epsilon_{2}}|\leq|\mathcal{N}_{\epsilon_{1}}(\Delta^{K})||\mathcal{N}_{\epsilon_{2}}|^{K}\leq\left(5\over\epsilon_{1}\right)^{K-1}\left(3\delta_{0}\over\overline{\sigma}\epsilon_{2}\right)^{KL}.$ (93)
Since for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, there exists ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (91) – (92) such that for all $k\in[K]$,
$|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|\leq|\widehat{M}_{k}({\bm{\omega}}^{\prime})-M_{k}({\bm{\omega}}^{\prime})|+|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})|,$ (94)
we first bound the second term:
$\begin{split}&\ \left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})\right|\\ &=\left|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j}))\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right)\right|\\ &\leq\|\widehat{\pi}-\pi_{{\bm{\omega}}^{*}}\|_{1}\max_{j\in[p]}\left|{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right|\\ &\leq 2\max_{j\in[p]}\left(|\alpha_{k}-\alpha_{k}^{\prime}|{A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\\ &\leq 2\max_{j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\\ &\leq 2\max_{j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right).\end{split}$ (95)
@@
Hence, invoking Lemma 7 gives
@@
$\quad\lesssim\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}.$ (97)
@@
[…]
@@
$\epsilon_{1}=\underline{\alpha}\left({1\over 4}\wedge{KL\over N}\right),\qquad\epsilon_{2}={1\over\overline{\sigma}}\left(\delta_{0}\wedge{\overline{\sigma}\over\|\mathbf{X}\|_{\infty,2}}{KL\over N}\right),\qquad t=CKL\log(N)$
and using (25) to collect terms.
@@
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}\leq\max_{{\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})\|_{2}$
@@
Pick any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ and ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (91). We bound from above
@@
where $\mathcal{N}_{L}(1/2)$ is the (1/2)(1/2)( 1 / 2 )-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/2)|\leq 5^{L}$ (see, for instance, Vershynin (2018)). Fix any ${\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ and $v\in\mathcal{N}_{L}(1/2)$. Observe that
@@
$\leq\alpha_{k}\max_{j\in[p]}\left(1+{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{*}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\lambda_{1}\left(\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}_{k}}}\mathbf{X}\right)$
$\overset{(i)}{\lesssim}\alpha_{k}\left(1+\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}+\delta_{0}\right)\lambda_{1}\left(H_{{\bm{\theta}}_{k}}\right)$
$\overset{(ii)}{\lesssim}\alpha_{k}~\overline{\sigma}^{2}$
[…]